OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

the line eq. will be y=3x+c where c is a constant

OpenStudy (anonymous):

to find c sub. by point (4,-2) in the equation

OpenStudy (anonymous):

what you mean

OpenStudy (anonymous):

the two lines are parallel then they have the same slopes (3)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

-2=12+c c=-14 the equation y=3x-14

OpenStudy (anonymous):

can you explain how did you set that up

OpenStudy (anonymous):

please

OpenStudy (anonymous):

first for create line equation by using the slope and a point that is the rule y=mx+c

OpenStudy (anonymous):

m=slope c=y-intercept

OpenStudy (anonymous):

now you have a point(4,-2) and you have a line parallel to the required line

OpenStudy (anonymous):

then the slope of the two lines is similar

OpenStudy (anonymous):

the slope of the line =3

OpenStudy (anonymous):

now you have a point =(4,-2) and m=3 use the rule (y-(-2))=m(x-4)

OpenStudy (anonymous):

y-2=3x-12 y=3x-14